OpenStudy (linn99123):

NEED HELP FAST :O I will fan and give medal Dawnie deposited $350 in a savings account earning 2% interest over 3 years. What was the total amount of interest earned at the end of 3 years? A. $420 B. $210 C. $42 D. $21

OpenStudy (anonymous):

D :*

OpenStudy (anonymous):

HEY I TECHNICALLY JUST GAVE YOU THE D

OpenStudy (the_fizicx99):

350 * 0.02 = 7 7 * 3 = 21

OpenStudy (anonymous):

f(x)=350*1.02^x

OpenStudy (anonymous):

x is the amount of years

OpenStudy (the_fizicx99):

That would be 371.4228

OpenStudy (anonymous):

371$

OpenStudy (the_fizicx99):

The formula actually is f(x) = P(r + 1)^x 371 isn't one of the options.

OpenStudy (anonymous):

371-350=21

OpenStudy (anonymous):

i didn't read the question :)